alones (alones):

SOMEONE CHECK !!!! Thanks :3

alones (alones):

(x^2 + 7x + 12) ÷ (x + 4) It's a long division

alones (alones):

x + 4 is my answer :3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\[x^2+7x+12=(x+3)(x+4)\] cancel

alones (alones):

Oh okay so it's wrong :3

alones (alones):

I had to write is as along division

OpenStudy (misty1212):

\[\frac{(x+3)(x+4)}{x+4}=?\]

rebeccaxhawaii (rebeccaxhawaii):